# Multiscale Latent Diffusion Model for Enhanced Feature Extraction from Medical Images

### v1

@@ Abstract
Several imaging modalities are utilized in the diagnosis of patients, each offering unique advantages and providing crucial insights into different aspects of anatomy and pathology. Among them, Computed tomography (CT) is a vital imaging modality in medical diagnostics, offering detailed cross-sectional images that allow clinicians to visualize internal structures with high precision. Its ability to detect subtle variations in tissue density makes it indispensable for diagnosing conditions such as lung cancer, where early detection and characterization of tumors are critical. However, inconsistencies in CT scanner types and acquisition protocols introduce significant variability in the extracted radiomic features, even when images are taken from the same patient. This variation complicates downstream research and clinical analysis, which depend on consistent feature representations. Current approaches for medical image features extraction often rely on supervised learning techniques, such as GAN-based models, but are limited in their ability to generalize across diverse imaging settings. We introduce LTDiff++, a multiscale latent diffusion model designed to extract enhanced features for medical images by transforming non-standard distributions into a standardized form within the latent space. The model employs a UNet++ encoder-decoder architecture, coupled with a conditional Denoising Diffusion Probabilistic Model (DDPM) at the latent bottleneck, to achieve robust feature extraction and standardization. The training process involves independently optimizing the encoder-decoder before integrating the latent DDPM for feature transformation. Empirical evaluations on both patient and phantom CT datasets demonstrate significant improvements in image standardization, achieving higher Concordance Correlation Coefficients (CCC) across multiple radiomic feature categories. Additionally, LTDiff++ effectively reduces image noise, confirming its potential for advanced medical imaging tasks.
@@
CT imaging, standardization, latent diffusion model, multiscale modeling, ensemble methods
@@ I Introduction
One of the most prevalent malignancies affecting both men and women in the US, lung cancer continues to be the leading cause of cancer-related mortality [1]. Non-small cell lung cancer (NSCLC) has a five-year survival rate of about 19%. In order to improve treatment outcomes, computed tomography (CT) imaging is crucial for the early diagnosis of lung cancer and for defining the features of the tumor [2, 3]. Additionally, texture analysis of CT images makes it easier to quantify temporal and geographical variations in the functionality and structure of tumors, which allows for the evaluation of intra-tumor evolution [4],[5]
@@ II-A CT Image Acquisition and Reconstruction Parameters
Figure 1: Shows how disparities in imaging procedures might result in differences in tumor picture characteristics. The same scanner was used to scan the identical lungman chest phantom. Two distinct image reconstruction techniques were utilized to get CT images.kernels appropriately, as the text at the bottom of the pictures shows.Green rectangles are used to indicate tumors in the images on the left (the top row shows the tumor regions that have been zoomed in). The feature variance in terms of CCC between these two tumors was displayed by the histogram on the right. The potential of extensive radiomic research may be significantly impacted by the detected variations in the tumor pictures.
@@ IV Results
Using four Nvidia V100 GPU cards on a Linux computer server, LTDiff++ was developed with PyTorch. Initialization of the network weights was achieved by tuning the parameters. The Adam optimizer was used to set the learning rate to 10-4. The encoder-decoder network was trained for 400 epochs. We trained the diffusion model with the same parameters and epochs. It took the model roughly 72 hours to train from the beginning. After training, the model processed and synthesized a DICOM CT image slice in approximately 30 seconds. We have evaluated the results of the model with concordance Correlation Coefficient (CCC) metric and relative error and downstream analysis for future lung cancer risk prediction.In addition to the original DDPM and the encoder-decoder network, we compared LTDiff++ with the current established method of standardization, DiffusionCT[11].
@@ V Ablation study
We conducted an ablation study to assess the impact of various components within the LTDiff++ framework on both phantom and patient data. Initially, we employed the conditional denoising diffusion probabilistic model (DDPM) to train the data, followed by an evaluation of the radiomic features. Subsequently, we compared the outcomes using the proposed LTDiff++ architecture, which was trained solely using the UNet++ encoder-decoder. Additionally, the training results of LTDiff++ model were analyzed by leveraging the skip connections of the UNet++ model configured as an encoder-decoder. As depicted in Table I, the images generated by LTDiff++ exhibited the highest Concordance Correlation Coefficient (CCC) values for the Gray Level Co-occurrence Matrix (GLCM) compared to those generated by other models, demonstrating a visual similarity to the standard image that was notably superior.

### v3

@@ Abstract
Various imaging modalities are used in patient diagnosis, each offering unique advantages and valuable insights into anatomy and pathology. Computed Tomography (CT) is crucial in diagnostics, providing high-resolution images for precise internal organ visualization. CT’s ability to detect subtle tissue variations is vital for diagnosing diseases like lung cancer, enabling early detection and accurate tumor assessment. However, variations in CT scanner models and acquisition protocols introduce significant variability in the extracted radiomic features, even when imaging the same patient. This variability poses considerable challenges for downstream research and clinical analysis, which depend on consistent and reliable feature extraction. Current methods for medical image feature extraction, often based on supervised learning approaches, including GAN-based models, face limitations in generalizing across different imaging environments. In response to these challenges, we propose LTDiff++, a multiscale latent diffusion model designed to enhance feature extraction in medical imaging. The model addresses variability by standardizing non-uniform distributions in the latent space, improving feature consistency. LTDiff++ utilizes a UNet++ encoder-decoder architecture coupled with a conditional Denoising Diffusion Probabilistic Model (DDPM) at the latent bottleneck to achieve robust feature extraction and standardization. Extensive empirical evaluations on both patient and phantom CT datasets demonstrate significant improvements in image standardization, with higher Concordance Correlation Coefficients (CCC) across multiple radiomic feature categories. Through these advancements, LTDiff++ represents a promising solution for overcoming the inherent variability in medical imaging data, offering improved reliability and accuracy in feature extraction processes.
@@
CT imaging, standardization, latent diffusion model, multiscale modeling
@@ I Introduction
One of the most prevalent malignancies affecting both men and women in the US, lung cancer continues to be the leading cause of cancer-related mortality [1]. Non-small cell lung cancer (NSCLC) has a five-year survival rate of about 19%. In order to improve treatment outcomes, computed tomography (CT) imaging is crucial for the early diagnosis of lung cancer and for defining the features of the tumor [2, 3]. Additionally, texture analysis of CT images makes it easier to quantify temporal and geographical variations in the functionality and structure of tumors, which allows for the evaluation of intra-tumor evolution [4],[5]. Image analysis technology is utilized to extract clinically relevant information from medical images, while image processing techniques enable the efficient and accurate segmentation of key features. Together, these technologies form an essential foundation for physicians in developing consultation strategies, surgical interventions, and diagnostic plans [6]. Moreover, due to the complexity, variability, and individual differences in organ structures within the human body, effectively extracting relevant features from medical images is critical for enhancing the accuracy of disease diagnosis and treatment planning. Currently, the most widely utilized techniques include the Gaussian curve method, direction template method, matrix method, and threshold method. Some researchers have introduced a novel approach for subpixel boundary extraction using Gaussian fitting. This approach involves selecting a set of points along the boundary of the original image, applying grayscale processing to those points, and then performing high-precision fitting for accurate boundary extraction [7], [8].
Feature extraction from CT images plays a pivotal role in medical imaging tasks, including segmentation, classification, and disease diagnosis. Traditional methods for feature extraction often relied on handcrafted techniques, but deep learning algorithms have revolutionized this field by automatically learning and extracting hierarchical features from raw images [9]. Convolutional Neural Networks (CNNs) have become the backbone of many medical imaging applications due to their ability to capture spatial hierarchies in pixel data. In particular, CNNs can be fine-tuned to recognize specific patterns, such as lesions or tumors, which may be critical for early diagnosis and treatment planning [10]. These models excel at capturing local features, such as edges and textures, while progressively learning more abstract representations in deeper layers.
Recently, diffusion models have emerged as a powerful generative approach for enhancing image quality and extracting subtle features that might be overlooked by traditional CNNs. Diffusion models are based on iterative denoising processes, where a noisy image is progressively refined to recover detailed structures [11]. This technique is particularly useful in medical imaging, where noise and low contrast often obscure important features in CT scans. For example, Denoising Diffusion Probabilistic Models (DDPMs) have been shown to effectively reconstruct high-quality images from noisy inputs, preserving fine details such as the boundaries of tumors or lesions [12]. By enhancing the visibility of these critical regions, diffusion models provide more informative features for downstream tasks, such as classification and segmentation.
Combining deep learning approaches with diffusion models offers a promising direction for medical imaging. CNNs provide a solid framework for feature extraction, while diffusion models augment the process by refining and enhancing these features, especially in challenging cases where noise or artifacts may degrade image quality. In particular, diffusion models can improve the clarity of regions of interest (ROI), such as tumors, which are often difficult to delineate in raw CT images. This fusion of techniques enables a more comprehensive analysis of the medical data, improving both the accuracy of diagnostic models and the interpretability of results. By leveraging the strengths of both deep learning and diffusion models, researchers are advancing the state-of-the-art in medical image analysis, providing more robust tools for clinical applications [13]. In comparison to established generative models such as Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs), Denoising Diffusion Probabilistic Models (DDPMs) have demonstrated superior performance in image standardization tasks. DDPMs use a Markov chain process that progressively transforms a simple starting distribution, like isotropic Gaussian noise, into the target data distribution. This approach involves two key stages: first, a forward diffusion process gradually adds noise to the image by sequentially sampling latent variables, and second, a reverse process uses a neural network, often a U-Net, to progressively remove the noise, recovering a clear image from the noisy input. Since gaining attention in 2020, DDPMs and their variants have driven major advancements in data modeling, with notable successes in areas like image generation, super-resolution, and image-to-image translation. More recently, conditional DDPMs have shown exceptional results in generating images based on specific conditions, while latent DDPMs have enabled efficient image generation within a low-dimensional latent space.
The training process for our proposed model, LTDiff++ involves three main phases. First, the multiscale UNet++ encoder-decoder network is trained using all available CT images, whether standardized or not, to learn how to encode each image into a latent vector that can be used to reconstruct the original image with minimal information loss. Next, a latent conditional DDPM is trained on pairs of non-standard and standardized images, allowing the model to learn the conditional probability distribution required to synthesize standardized images. Finally, the trained networks are combined to get the enhanced features from new CT images, enabling the robust harmonization of image data while maintaining crucial structural features.
@@ II-A CT Image Acquisition and Reconstruction Parameters
Figure 1: Shows how disparities in imaging procedures might result in differences in tumor picture characteristics. The same scanner was used to scan the identical lungman chest phantom. Two distinct image reconstruction techniques were utilized to get CT images kernels appropriately, as the text at the bottom of the pictures shows. Red squares are used to indicate tumors in the images on the left(tumor zoomed in the top left corner). The feature variance in terms of CCC between these two tumors was displayed by the histogram on the right. The potential of extensive radiomic features may be significantly impacted by the detected variations in the tumor pictures.
@@ IV Results
Using four Nvidia V100 GPU cards on a Linux computer server, LTDiff++ was developed with PyTorch. Initialization of the network weights was achieved by tuning the parameters. The Adam optimizer was used to set the learning rate to 10-4. The encoder-decoder network was trained for 400 epochs. We trained the diffusion model with the same parameters and epochs. It took the model roughly 72 hours to train from the beginning. After training, the model processed and synthesized a DICOM CT image slice in approximately 30 seconds. We have evaluated the results of the model with concordance Correlation Coefficient (CCC) metric and relative error and downstream analysis for future lung cancer risk prediction.In addition to the original DDPM and the encoder-decoder network, we compared LTDiff++ with the current established method of standardization, DiffusionCT[19]. Figure 3 illustrates the outcomes of the models on a sample tumor, highlighting noticeable differences between the input tumor image and the standard image, both in visual appearance and radiomic characteristics. The image produced by LTDiff++ stands out, showing the highest Concordance Correlation Coefficient (CCC) values for Gray-Level Co-occurrence Matrix (GLCM) features when compared to the standard image.
@@ V Ablation study
We conducted an ablation study to assess the impact of various components within the LTDiff++ framework on both phantom and patient data. Initially, we employed the conditional denoising diffusion probabilistic model (DDPM) to train the data, followed by an evaluation of the radiomic features. Subsequently, we compared the outcomes using the proposed LTDiff++ architecture, which was trained solely using the UNet encoder-decoder and with different loss like L1 loss.From Table I we can see that LTDiff++ consistently outperforms the Baseline, UNet++, and DiffusionCT models across all radiomic feature groups, achieving near-perfect CCC values for GOH (1.00 ± 0.00) and substantial improvements in features such as GLCM (0.86 ± 0.11), GLRLM (0.84 ± 0.04), IH (0.72 ± 0.04), and NID (0.88 ± 0.03). This indicates that LTDiff++ effectively preserves radiomic features during image synthesis. In contrast, the ablation study results in Table III evaluate the effect of modifying LTDiff++ by either incorporating UNet or replacing it with L1 loss. The results show that LTDiff++ with UNet achieves superior performance (e.g., GOH: 75.00 ± 3.4 and GLCM: 0.71 ± 0.35), indicating the importance of UNet in maintaining feature integrity. However, replacing UNet with L1 loss results in a significant performance drop (e.g., GOH: 55.00 ± 2.4 and GLCM: 0.61 ± 0.35), demonstrating the critical role of UNet in the model’s architecture. When compared to Table I, the ablation study confirms that the full LTDiff++ model integrates its components effectively to achieve state-of-the-art performance, while modifications to its structure lead to reduced concordance with original patient data.
TABLE III: Ablation study for Patient data
Model GOH GLCM GLRLM ID IH NID
LTDiff++ version_1 with UNet 75.00 $\pm$ .34 0.71 $\pm$ 0.35 0.62 $\pm$ 0.26 0.78 $\pm$ 0.45 0.58 $\pm$ 0.14 0.57 $\pm$ 0.56
LTDiff++ version_1 with L1 loss 55.00 $\pm$ .24 0.61 $\pm$ 0.35 0.56 $\pm$ 0.16 0.68 $\pm$ 0.25 0.52 $\pm$ 0.24 0.54 $\pm$ 0.19